# Echoes and defects in the Calogero model

### v1

@@ Abstract
In this work, we extend the study of the interplay between scaling symmetries and statistics to one-dimensional fluids by studying the Calogero model in a harmonic trap modulated through time. The latter harbors an interpretation in terms of free particles imbued with exclusion statistics and is an example of a scale invariant fluid in 1D displaying $SO(2,1)$ hidden symmetry preserved by harmonic traps. Taking advantage of the dynamical symmetry, two experimentally-relevant drive protocols spanning both quasi-static and non-adiabatic regimes are investigated and universal signatures of the interactions and exclusion statistics are uncovered in the ground state echo amplitude and closely related ground state fidelity. In particular, under both periodic modulation and slow drive through the gapless point of the trap frequency, enhanced interactions and exclusion are shown to favor the proliferation of defects and to hinder their annihilation, which leads to a universal decrease of ground state fidelities and echo amplitudes. We also show that increasing exclusion sparks a sharp suppression of the likelihood of intermediate echoes beyond those imposed by the commensurability of a periodic drive and the natural frequency of the trap.
@@ I Introduction
Symmetries are fundamental to our understanding of physical systems. The associated existence of conserved quantities greatly simplifies the study of their dynamics. More generally, a system might harbor dynamical symmetries (also referred to as hidden or spectrum-generating symmetries) resulting in additional dynamical invariants. In particular, scale invariant fluids form a notable class of systems that displays rich dynamical symmetries. A defining feature of such fluids is their behavior under a dilation $x\to\zeta x$, for which the Hamiltonian scales as $H(\zeta x)=\zeta^{\alpha}H(x)$ where $\alpha$ is some real exponent. Such scaling has dramatic consequences: for example, it leads to exceptionally simple thermodynamic properties such as equations of state depending solely on the ratio of temperature to chemical potential as opposed to both independently Saint-Jalm et al. (2019). Remarkably, even when confined by a parabolic potential, scale invariant fluids may still display $SO(2,1)$ hidden symmetry, again with deep consequences both for the time evolution and equilibrium properties of the fluid.
Experimental observations of effectively scale invariant fluids are ubiquitous, notably thanks to ultracold atomic gas platforms. This includes a myriad of systems among which the unitary Fermi gas in 3D Castin and Werner (2012), the 2D Bose gas Pitaevskii and Rosch (1997) in the weakly interacting limit and 2D Fermi gas away from the crossover regime Hofmann (2012); Murthy et al. (2019), the Tonks-Girardeau gas in 1D Minguzzi and Gangardt (2005) as well as all systems effectively described through a Gross-Pitaevskii equation in all dimensions Pérez-García et al. (1997).
In this context, the $SO(2,1)$ dynamical symmetry Lv et al. (2020) was shown to be responsible for the emergence of breathers and echoes in two-dimensional Bose gas experiments Saint-Jalm et al. (2019). These studies focused on the dynamics of $SO(2,1)$-symmetric systems prepared in the ground state of their non-interacting counterparts (i.e. the ground state of the trapping Hamiltonian). The crucial presence of a $SO(2,1)$ dynamical symmetry equally enables the study of the dynamics initialized in the fully interacting ground state, which is the focus of the present work.
@@
Lower dimensional many-body systems offer the potential for generalized notions of statistics, beyond the bosonic and fermionic limits, raising the interest for the exploration of the interplay between scaling symmetries and statistics in this setting. In this work, we consider one dimension, where a different notion of statistics is often considered: exclusion statistics Haldane (1991). The Calogero model Calogero (1971); Sutherland (1971, 1972); Moser (1975) is a paradigmatic one-dimensional model displaying both a scaling symmetry and an interpretation in terms of free particles imbued with exclusion statistics Ujino and Wadati (1995). This model displays the dynamical $SO(2,1)$ symmetry alluded to above, making it an ideal system to explore the interplay between exclusion statistics and hidden symmetries. In this work, we investigate the dynamics of the Calogero model in a time-dependent harmonic trap, which explicitly reduces the conformal invariance to the $SO(2,1)$ spectrum generating symmetry. We study both periodic modulations of the trap frequency and slow drives through the gapless point. Enhanced exclusion and interactions are predicted to favor the proliferation of defects and to hinder their annihilation. Our claims are then verified through the computation of ground state fidelities and echo amplitudes, in which stronger exclusion and interactions are shown to universally lower the chance of revivals.
@@ II Model, symmetries, and exclusion interpretation
The eigenfunctions $\Psi_{n}(\left\{x_{i}\right\})$ of the Calogero model are known exactly Vacek et al. (1994). In particular, the ground state takes the form
@@
This can be interpreted as an exclusion principle that requires the pseudo-excitation numbers to be at least $\lambda$ apart, which manifests a generalized statistics interpretation Ujino and Wadati (1995). Particular cases include the free boson gas at $\lambda=0$ and the Tonks-Girardeau gas (hard-core bosons) at $\lambda=1$, while other values describe one dimensional Haldane anyons Haldane (1991). Note that the numbers $\bar{n}_{i}$ are not necessarily integers, but they are modified through integer increments. The minimal allowed nonnegative set of values for $\bar{n}_{i}$ obeying the selection rules determines the ground state uniquely while excited states are built through all integer increments thereof. Fig. 1 displays a schematic representation of this exclusion statistics interpretation.
Notably, the scaling of the kinetic energy and pair-wise interaction in (1) are identical under dilation, making the Calogero Hamiltonian in the absence of a trap scale-invariant. As pointed out earlier and as we will show in the following, the presence of a harmonic trapping potential breaks scale invariance but preserves the $SO(2,1)$ dynamical symmetry Pitaevskii and Rosch (1997), independently of whether the trapping potential is modulated through time, i.e. when $\omega=\omega(t)$. This allows us to go beyond static properties of the Calogero model and to uncover more general features of its dynamics associated with its $SO(2,1)$ hidden symmetry. A direct consequence of the dynamical symmetry is the predictable dynamics of the time-dependent many-body wavefunction $\psi_{n}(t)$ as the harmonic frequency is varied in time $\omega(t)$. We indeed have that if the system occupies the eigenstate $\Psi_{n}$ with trapping potential $\omega(t_{0})=\bar{\omega}$ at any instant $t_{0}$, the time-evolved state at time $t$ with trapping frequency $\omega(t)$ can be obtained through the formula del Campo (2011); Gritsev et al. (2010)
@@
We now show how scale invariance can be argued to lead to a $SO(2,1)$ dynamical symmetry even in the presence of a time-dependent harmonic trap, and sketch how scaling symmetry leads to the scaling dynamics described by Eq. (5). Let us assume a Hamiltonian of the form ${H}(t)=H_{0}+\omega^{2}(t)H_{\mathrm{trap}}$, with $H_{0}=K+V$ where $K=-\tfrac{1}{2}\sum_{i}\partial_{i}^{2}$ is the standard non-relativistic kinetic Hamiltonian, $H_{\mathrm{trap}}=\tfrac{1}{2}\sum_{i}x_{i}^{2}$ the unit-frequency harmonic trap and the potential $V$ scales like the kinetic part $K$ under rescaling $V(\{\zeta x_{i}\})=\zeta^{-2}V(\{x_{i}\})$. This form encompasses in particular the driven Calogero model. As shown in Gritsev et al. (2010), taking Eq. (5) as an ansatz for the dynamical state, imposing $i\dot{\psi}=H(t)\psi$ and using the assumed scaling of the potential $V$ straightforwardly results in $\xi$ satisfying the Ermakov equation (6). The occurrence of these scaling dynamics is intimately linked to the fact that $V$ scales like $K$ under rescaling, or in other words that the untrapped Hamiltonian $H_{0}$ is scale invariant. A special property of the harmonic potential $H_{\mathrm{trap}}$ is the fact that its commutator with any scale-invariant base Hamitonian yields the generator of scaling transformations: $[H_{\mathrm{trap}},H_{0}]=iQ$ where $Q=\tfrac{1}{2}\sum_{j}\left((-i\partial_{j})x_{j}+x_{j}(-i\partial_{j})\right)$. Consequently, a closed $SO(2,1)$ algebra can be generated out of linear combinations of $H_{\mathrm{trap}},H_{0}$ and $Q$. Defining $T_{1}=\frac{1}{2}\left(H_{0}-H_{\mathrm{trap}}\right)$, $T_{2}=\frac{Q}{2}$ and $T_{3}=\frac{1}{2}\left(H_{0}+H_{\mathrm{trap}}\right)$, we indeed find that $\left[T_{1},T_{2}\right]=-iT_{3},~{}~{}\left[T_{2},T_{3}\right]=iT_{1},~{}~{}\left[T_{3},T_{1}\right]=iT_{2}$. The time-dependent Hamiltonian can then be explicitly expressed in terms of these generators as $H(t)=(1-\omega(t)^{2})T_{1}+(1+\omega(t)^{2})T_{3}$. The associated evolution operator can subsequently be factorized into a product of exponentials of the generators $T_{i}$ Martínez-Tibaduiza et al. (2020); Ban (1993). The ”spectrum-generating” action of these generators transforming eigenstates into one another is known and thereby formally solves the full time evolution, resulting in Eq. (5).
All in all, we have argued that given a solution of the Ermakov-Milne equation for a particular functional form of the modulated frequency $\omega(t)$, Eq. (5) provides the exact evolution of the state of the system when prepared in one of its eigenstates. In this paper, we study the system prepared in its ground state $\Psi_{0}$. Particularizing (5) to $n=0$ and using the expression for $\Psi_{0}$ and $E_{0}$ following from (2) and (3), we straightforwardly find an expression for the evolved ground state of the Calogero model $\psi_{0}(t)$ in terms of the solution $\xi(t)$ of the Ermakov-Milne equation
@@
Figure 1: Schematic representation of the exclusion statistics interpretation of the Calogero model. The central black box represents the evenly spaced many-body spectrum with an offset by $N(N-1)\lambda\omega/2$ and each state is labeled by $n$. Each such many-body state $n$ can be viewed as resulting from the filling of a single-particle spectrum without offset with $N$ particles labeled $i$ together with the constraint that the single-particle level $\bar{n}_{i}$ occupied by the $i$th particle is at least separated by $\lambda$ from the level of the previous particle $\bar{n}_{i-1}$. An illustration is given for $\lambda=2$ and $N=4$. The many-body ground state in grey in the central black box corresponds to the configuration minimizing the numbers $\bar{n}_{i}$ as shown in the red box on the left. The first excited state corresponds to a configuration which increases these numbers by one unit as shown in the blue box on the right. And so on for higher excited states.
@@ III Defect proliferation out of equilibrium
We first present a heuristic argument for how statistics affect defect proliferation when the harmonic trap frequency $\omega$ is varied in time, as illustrated in Fig. 2. Note that when $\omega$ is lowered, the energy level spacing decreases and the spectrum has a downward shift in energy proportional to the interaction strength $\lambda$. Starting in the ground state $E_{0}$ at some fixed $\omega$, as $\omega$ decreases, the number of energy levels with higher quantum number $n$ that cross (below) the initial energy level $E_{0}$ grows with $\lambda$. Large values of $\lambda$ are therefore expected to favor transitions to higher quantum numbers when the trapping potential is weakened. Interactions thus facilitate the proliferation of defects. Conversely, when $\omega$ increases, the spectrum is shifted up and the level spacing increases. The energy of the populated levels with highest quantum numbers increases fast with respect to the energy of the ground state, which renders the annihilation of defects more difficult.
The physical picture provided above essentially stems from an exclusion constraint on free particles. In the single-particle energy spectrum picture, there is no longer a $\lambda$-dependent shift but rather the constraint for the quantum numbers $\bar{n}_{i}$ to be at least $\lambda$ apart. As the spectrum is filled from the bottom with a fixed number of particles $N$, the single-particle energies occupied are more spread out and larger for increasing $\lambda$. When the level spacing is lowered through a modulation of $\omega$, the increased density of states and the greater initial span in energy of particles lower the importance of the constraints. This offers more reconfiguration possibilities into single-particle states with higher quantum numbers $\bar{n}_{i}$. As the level spacing is restored, not only are the defects to be annihilated at higher levels, the constraints become more restricting and the number of relaxation paths from these higher states to the ground state is reduced.
In our analysis, we focus on two closely related observables, relevant in cold atom experiments. In particular, we wish to investigate in both cases the deviation of the dynamical state (harboring defects) with respect to a given choice of reference state. For quasi-static drives, we focus on the ground state fidelity $f(t)$ with the instantaneous adiabatic ground state, while in non-adiabatic settings, as the initial state is more relevant we consider the closely related echo amplitude $e(t)$. These are defined as
@@
where $\psi^{ad}_{0}(t)$ is the instantaneous equilibrium ground state, $\psi_{0}(0)$ is the initial ground state and $\psi_{0}({t})$ is the time-evolved dynamical state. Based on the heuristic discussion presented earlier, we expect the enhanced proliferation of defects sparked by higher values of $\lambda$ to induce a decrease in the ground state echo amplitude and fidelity.
Figure 2: Schematic representation of the experimental protocols investigated in this work. Starting with the system in its ground state (in grey in the black box on the left) trapped with a given frequency $\omega=\omega_{+}$, the first protocol drives the frequency $\omega$ through 00 (third box in blue) before restoring the trapping frequency $\omega=\omega_{+}$ to its original value. The second protocol (second box in red) is designed in a similar fashion except for the fact that the initial trapping frequency $\omega_{+}$ is driven to a more general lower value $\omega_{-}$ before being restored and that this procedure is repeated with a period $2\pi/\Omega$. In both cases, while modulating the trapping frequency $\omega$, higher values of $\lambda$ favor transitions to states with higher quantum numbers and hinder the return to the ground state (green arrows) and therefore facilitate the proliferation of defects.

### v2

@@ Abstract
In this work, we extend the study of the interplay between scaling symmetries and statistics to one-dimensional fluids by studying the Calogero model in a harmonic trap modulated through time. The latter harbors an interpretation in terms of free particles imbued with exclusion statistics and is an example of a scale invariant fluid in one-dimension displaying SO(2,1) dynamical symmetry preserved by harmonic traps. Taking advantage of the dynamical symmetry, two experimentally relevant drive protocols spanning both quasi-static and non-adiabatic regimes are investigated and universal signatures of the interactions and exclusion statistics are uncovered in the ground-state echo amplitude and closely related ground-state fidelity. In particular, under both periodic modulation and slow drive through the gapless point of the trap frequency, enhanced interactions and exclusion are shown to favor the proliferation of defects and to hinder their annihilation, which leads to a universal decrease of ground-state fidelities and echo amplitudes. We also show that increasing exclusion sparks a sharp suppression of the likelihood of intermediate echoes beyond those imposed by the commensurability of a periodic drive and the natural frequency of the trap.
@@ I Introduction
Symmetries are fundamental to our understanding of physical systems. The associated existence of conserved quantities greatly simplifies the study of their dynamics. More generally, a system might harbor dynamical symmetries (also referred to as hidden or spectrum-generating symmetries) resulting in additional dynamical invariants. In particular, scale invariant fluids form a notable class of systems that displays rich dynamical symmetries. A defining feature of such fluids is their behavior under a dilation $x\to\zeta x$, for which the Hamiltonian scales as $H(\zeta x)=\zeta^{\alpha}H(x)$ where $\alpha$ is some real exponent. Such scaling has dramatic consequences: for example, it leads to exceptionally simple thermodynamic properties such as equations of state depending solely on the ratio of temperature to chemical potential as opposed to both independently Saint-Jalm et al. (2019). Remarkably, even when confined by a parabolic potential, scale invariant fluids may still display $\mathrm{SO(2,1)}$ hidden symmetry, again with deep consequences both for the time evolution and equilibrium properties of the fluid.
Experimental observations of effectively scale invariant fluids are ubiquitous, notably thanks to ultracold atomic gas platforms. This includes a myriad of systems among which the unitary Fermi gas in three dimensions Castin and Werner (2012), the two-dimensional (2D) Bose gas Pitaevskii and Rosch (1997) in the weakly interacting limit and 2D Fermi gas away from the crossover regime Hofmann (2012); Murthy et al. (2019), the Tonks-Girardeau gas in one dimension Minguzzi and Gangardt (2005) as well as all systems effectively described through a Gross-Pitaevskii equation in all dimensions Pérez-García et al. (1997).
In this context, the $\mathrm{SO(2,1)}$ dynamical symmetry Lv et al. (2020) was shown to be responsible for the emergence of breathers and echoes Lyu et al. (2020) in 2D Bose gas experiments Saint-Jalm et al. (2019). These studies focused on the dynamics of $\mathrm{SO(2,1)}$-symmetric systems prepared in the ground-state of their non-interacting counterparts (i.e. the ground-state of the trapping Hamiltonian). The crucial presence of a $\mathrm{SO(2,1)}$ dynamical symmetry equally enables the study of the dynamics initialized in the fully interacting ground-state, which is the focus of the present work.
@@
Lower dimensional many-body systems offer the potential for generalized notions of statistics, beyond the bosonic and fermionic limits, raising the interest for the exploration of the interplay between scaling symmetries and statistics in this setting. In this work, we consider one dimension, where a different notion of statistics is often considered: exclusion statistics Haldane (1991). The Calogero model Calogero (1971); Sutherland (1971, 1972); Moser (1975) is a paradigmatic one-dimensional model displaying both a scaling symmetry and an interpretation in terms of free particles imbued with exclusion statistics Ujino and Wadati (1995). This model displays the dynamical $\mathrm{SO(2,1)}$ symmetry alluded to above, making it an ideal system to explore the interplay between exclusion statistics and hidden symmetries. In this work, we investigate the dynamics of the Calogero model in a time-dependent harmonic trap, which explicitly reduces the conformal invariance to the $\mathrm{SO(2,1)}$ spectrum generating symmetry. We study both periodic modulations of the trap frequency and slow drives through the gapless point. Enhanced exclusion and interactions are predicted to favor the proliferation of defects and to hinder their annihilation. Our claims are then verified through the computation of ground-state fidelities and echo amplitudes, in which stronger exclusion and interactions are shown to universally lower the chance of revivals.
@@ II Model, symmetries, and exclusion interpretation
The eigenfunctions $\Psi_{n}(\left\{x_{i}\right\})$ of the Calogero model are known exactly Vacek et al. (1994). In particular, the ground-state takes the form
@@
This can be interpreted as an exclusion principle that requires the pseudo-excitation numbers to be at least $\lambda$ apart, which manifests a generalized statistics interpretation Ujino and Wadati (1995). Particular cases include the free boson gas at $\lambda=0$ and the Tonks-Girardeau gas (hard-core bosons) at $\lambda=1$, while other values describe one dimensional Haldane anyons Haldane (1991). Note that the numbers $\bar{n}_{i}$ are not necessarily integers, but they are modified through integer increments. The minimal allowed nonnegative set of values for $\bar{n}_{i}$ obeying the selection rules determines the ground-state uniquely while excited states are built through all integer increments thereof. Figure 1 displays a schematic representation of this exclusion statistics interpretation.
Notably, the scaling of the kinetic energy and pair-wise interaction in Eq. (1) are identical under dilation, making the Calogero Hamiltonian in the absence of a trap scale-invariant. As pointed out earlier and as we will show in the following, the presence of a harmonic trapping potential breaks scale invariance but preserves the $\mathrm{SO(2,1)}$ dynamical symmetry Pitaevskii and Rosch (1997), independently of whether the trapping potential is modulated through time, i.e. when $\omega=\omega(t)$. This allows us to go beyond static properties of the Calogero model and to uncover more general features of its dynamics associated with its $\mathrm{SO(2,1)}$ hidden symmetry. A direct consequence of the dynamical symmetry is the predictable dynamics of the time-dependent many-body wavefunction $\psi_{n}(t)$ as the harmonic frequency is varied in time $\omega(t)$. We indeed have that if the system occupies the eigenstate $\Psi_{n}$ with trapping potential $\omega(t_{0})=\bar{\omega}$ at any instant $t_{0}$, the time-evolved state at time $t$ with trapping frequency $\omega(t)$ can be obtained through the formula del Campo (2011); Gritsev et al. (2010)
@@
We now show how scale invariance can be argued to lead to a $\mathrm{SO(2,1)}$ dynamical symmetry even in the presence of a time-dependent harmonic trap, and sketch how scaling symmetry leads to the scaling dynamics described by Eq. (5). Let us assume a Hamiltonian of the form ${H}(t)=H_{0}+\omega^{2}(t)H_{\mathrm{trap}}$, with $H_{0}=K+V$ where $K=-\tfrac{1}{2}\sum_{i}\partial_{i}^{2}$ is the standard non-relativistic kinetic Hamiltonian, $H_{\mathrm{trap}}=\tfrac{1}{2}\sum_{i}x_{i}^{2}$ the unit-frequency harmonic trap and the potential $V$ scales like the kinetic part $K$ under rescaling $V(\{\zeta x_{i}\})=\zeta^{-2}V(\{x_{i}\})$. This form encompasses in particular the driven Calogero model. As shown in Ref. Gritsev et al. (2010), taking Eq. (5) as an ansatz for the dynamical state, imposing $i\dot{\psi}=H(t)\psi$ and using the assumed scaling of the potential $V$ straightforwardly results in $\xi$ satisfying the Ermakov equation (6). The occurrence of these scaling dynamics is intimately linked to the fact that $V$ scales like $K$ under rescaling, or in other words that the untrapped Hamiltonian $H_{0}$ is scale invariant. A special property of the harmonic potential $H_{\mathrm{trap}}$ is the fact that its commutator with any scale-invariant base Hamitonian yields the generator of scaling transformations: $[H_{\mathrm{trap}},H_{0}]=iQ$ where $Q=\tfrac{1}{2}\sum_{j}\left[(-i\partial_{j})x_{j}+x_{j}(-i\partial_{j})\right]$. Consequently, a closed $\mathrm{SO(2,1)}$ algebra can be generated out of linear combinations of $H_{\mathrm{trap}},H_{0}$ and $Q$. Defining $T_{1}=\frac{1}{2}\left(H_{0}-H_{\mathrm{trap}}\right)$, $T_{2}=\frac{Q}{2}$ and $T_{3}=\frac{1}{2}\left(H_{0}+H_{\mathrm{trap}}\right)$, we indeed find that $\left[T_{1},T_{2}\right]=-iT_{3},~{}~{}\left[T_{2},T_{3}\right]=iT_{1},~{}~{}\left[T_{3},T_{1}\right]=iT_{2}$. The time-dependent Hamiltonian can then be explicitly expressed in terms of these generators as $H(t)=[1-\omega(t)^{2}]T_{1}+[1+\omega(t)^{2}]T_{3}$. The associated evolution operator can subsequently be factorized into a product of exponentials of the generators $T_{i}$ Martínez-Tibaduiza et al. (2020); Ban (1993). The ”spectrum-generating” action of these generators transforming eigenstates into one another is known and thereby formally solves the full time evolution, resulting in Eq. (5).
All in all, we have argued that given a solution of the Ermakov-Milne equation for a particular functional form of the modulated frequency $\omega(t)$, Eq. (5) provides the exact evolution of the state of the system when prepared in one of its eigenstates. In this paper, we study the system prepared in its ground-state $\Psi_{0}$. Particularizing Eq. (5) to $n=0$ and using the expression for $\Psi_{0}$ and $E_{0}$ following from Eqs. (2) and (3), we straightforwardly find an expression for the evolved ground-state of the Calogero model $\psi_{0}(t)$ in terms of the solution $\xi(t)$ of the Ermakov-Milne equation Sutherland (1998)
@@
Figure 1: Schematic representation of the exclusion statistics interpretation of the Calogero model. The central black box represents the evenly spaced many-body spectrum with an offset by $N(N-1)\lambda\omega/2$ and each state is labeled by $n$. Each such many-body state $n$ can be viewed as resulting from the filling of a single-particle spectrum without offset with $N$ particles labeled $i$ together with the constraint that the single-particle level $\bar{n}_{i}$ occupied by the $i$th particle is at least separated by $\lambda$ from the level of the previous particle $\bar{n}_{i-1}$. An illustration is given for $\lambda=2$ and $N=4$. The many-body ground-state in gray in the central black box corresponds to the configuration minimizing the numbers $\bar{n}_{i}$ as shown in the red box on the left. The first excited state corresponds to a configuration which increases these numbers by one unit as shown in the blue box on the right, and so on for higher excited states.
@@ III Defect proliferation out of equilibrium
Figure 2: Schematic representation of the protocol corresponding to an adiabatic drive through the gapless point. Starting with the system in its ground-state (in gray in the leftmost black box) trapped with a given frequency $\omega=\omega_{+}$, the first protocol continuously drives the frequency $\omega$ through 00 (third box in blue with full level degeneracy) before restoring the trapping frequency $\omega=\omega_{+}$ to its original value.
We first present a heuristic argument for how statistics affect defect proliferation when the harmonic trap frequency $\omega$ is varied in time. Figures 2 and 3, which depict specific choices of drive protocols for $\omega$ and their effects on the many-body spectrum, are both useful in illustrating the analysis.
Note that when $\omega$ is lowered, the energy level spacing decreases and the spectrum has a downward shift in energy proportional to the interaction strength $\lambda$. Starting in the ground-state $E_{0}$ at some fixed $\omega$, as $\omega$ decreases, the number of energy levels with higher quantum number $n$ that cross (below) the initial energy level $E_{0}$ grows with $\lambda$. Large values of $\lambda$ are therefore expected to favor transitions to higher quantum numbers when the trapping potential is weakened. Interactions thus facilitate the proliferation of defects. Conversely, when $\omega$ increases, the spectrum is shifted up and the level spacing increases. The energy of the populated levels with highest quantum numbers increases fast with respect to the energy of the ground-state, which renders the annihilation of defects more difficult.
The physical picture provided above essentially stems from an exclusion constraint on free particles. In the single-particle energy spectrum picture, there is no longer a $\lambda$-dependent shift but rather the constraint for the quantum numbers $\bar{n}_{i}$ to be at least $\lambda$ apart. As the spectrum is filled from the bottom with a fixed number of particles $N$, the single-particle energies occupied are more spread out and larger for increasing $\lambda$. When the level spacing is lowered through a modulation of $\omega$, the increased density of states and the greater initial span in energy of particles lower the importance of the constraints. This offers more reconfiguration possibilities into single-particle states with higher quantum numbers $\bar{n}_{i}$. As the level spacing is restored, not only are the defects to be annihilated at higher levels, but the constraints become more restricting and the number of relaxation paths from these higher states to the ground-state is reduced.
In our analysis, we focus on two closely related observables, relevant in cold atom experiments. In particular, we wish to investigate in both cases the deviation of the dynamical state (harboring defects) with respect to a given choice of reference state. For quasi-static drives, we focus on the ground-state fidelity $f(t)$ with the instantaneous adiabatic ground-state, while in non-adiabatic settings, as the initial state is more relevant we consider the closely related echo amplitude $e(t)$. These are defined as
@@
where $\psi^{ad}_{0}(t)$ is the instantaneous equilibrium ground-state, $\psi_{0}(0)$ is the initial ground-state and $\psi_{0}({t})$ is the time-evolved dynamical state. Based on the heuristic discussion presented earlier, we expect the enhanced proliferation of defects sparked by higher values of $\lambda$ to induce a decrease in the ground-state echo amplitude and fidelity.